DX LUCI AO Commissioning

07:54 Preparing to run the first script.

07:59
  • Preset AO Successful
  • Bin 2
  • 153 Modes
  • Frequency 990 Hz
  • Magnitude 9.8
  • Gopt Gain: 0.25
  • Gain: 0.46 0.20 0.14
  • Dataset 20170320_080254
  • Seeing 0.75-0.80
  • Tip-Tilt Voltage 7.0
08:08
  • Dataset 20170320_080922
  • Seeing went up to about 1.0.
08:20
  • Spot looks elongated on the AGW.
  • Dataset 20170320_082128
08:40
  • SX had loop open.
08:47
  • Beginning new preset.
  • DX failed because WFSARB was stopped. My fault.
08:58
  • New-New Preset.
  • Gains 1.97 0.81 0.76
  • 0.21 Gopt Gain.
  • Gopt on
  • Bin 2
  • 990 Hz
  • 153 Modes
  • GCS image still very elongated.
  • Dataset: 20170310_090942
09:15
  • We need to abort the preset on the Right.
  • Waiting for some offset problems to be resolved.
09:23
  • Preset.
  • Same Target.
  • RA: 13:38:53.2189548412757
  • DEC: +38:51:21.3030928254477
  • Bin 2
  • 153 Modes
  • 990 Hz
  • Gain 1.74 1.16 0.63
  • Gopt Gain 0.25
  • Dataset: 20170320_092943
09:33
  • Re-authorizing for only DX side.
  • New Preset
  • RA: 13:38:53.2189548412757
  • Dec: +38:51:21.3030928254477
  • Mag 10.0
  • Bin 1
  • 400 modes
  • 990 Hz
  • Gain 1.38 0.68 0.40
  • Gopt Gain 0.27
  • Reduced HO gain by 0.05 after GOPT Started
  • Gain 0.32 0.18 0.12
  • Dataset: 20170320_093952
  • Dataset: 20170320_094623
09:48
  • We offset to center the target in the fieldstop.
  • Continuing with same preset.
  • Same target coords in the data dictionary.
  • Dataset: 20170320_095045
  • Continuing smoothly
  • Dataset: 20170320_101303
  • Dataset: 20170320_103138
  • Several very large offsets were made.
  • Dataset: 20170320_104237, taken just after large offsets.
  • 0.48" on guider.
  • Dataset: 20170320_110058
11:10
  • DIMM 0.64
  • Dataset: 20170320_111101
  • Magnitude 9.9
11:37
  • New Preset
  • Seeing about 0.65".
  • RA: 14:06:47.7422933172926
  • DEC: +34:11:34.0666126067327
  • Bin 2
  • 625 Hz
  • Gain 1.70 1.10 0.62
  • Optical gain: 0.22
  • Dataset: 20170320_114218
  • Seeing rose to about 0.8
  • Dataset: 20170320_114851
11:55
  • Loop opened during offset
  • Optical gain may have become unstable.
  • New Preset
  • Gain 1.70 1.10 0.42
  • Leaving GOPT off as it may be slightly unstable at this loop speed.
  • Still Bin 2
  • 625 Hz
  • Magnitude 12.1
  • Dataset: 20170320_120548
  • Dataset: 20170320_121058
  • Dataset: 20170320_122216
  • Jochen complains about elongated images. This also is apparent in the GCS image. More study is needed.
  • Taking a bunch of datasets to see what things look like as it gets lighter.
  • Dataset: 20170320_123621
  • Dataset: 20170320_124414
  • Dataset: 20170320_124903
  • Dataset: 20170320_125219
12:54 Loop becoming unstable.

-- %USERSIG{GregoryTaylor - 2017-03-20}%
Topic revision: r2 - 22 Mar 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback